### v1

@@ 1 Introduction
Transfer based architectures, originally designed for natural language processing, have recently shown its promises in the field of image processing. Vision Transformers (ViTs), with their reliance on self-attention mechanisms, have demonstrated superior performance in tasks such as image recognition, segmentation, and multimodal learning [1]. As the medical domain is known for the scarcity of data, transfer-based architectures will also play an important role in this field in terms of segmentation, anomaly detection, and eventually image recognition, i.e., diagnosis. Unlike traditional Convolutional Neural Networks (CNNs) that process localized spatial features, ViTs operate on global relationships across image patches, allowing them to capture long-range dependencies effectively. This enables them to outperform CNNs in specific scenarios, including subtle pattern identification in complex medical images. However, this global attention mechanism also exposes them to unique vulnerabilities, particularly in adversarial settings.
@@
ViT take images, in this case - medical images, and divide them up into patches of a fixed size, and then maps these into embeddings linearly. This patch-based representation lacks the sort of hierarchical feature extraction that even the simplest CNNs perform, where lower layers can recognize simple shapes and higher-order ones recognize much more complex features. In contrast, the attention mechanism in ViT considers all patches simultaneously for the computation of their relationships. Therefore, scaling perturbations that change the interrelationships of patches can be used to cause misrepresentation of attention maps with respect to the original structure of the image. ViTs also operate on random global attention mechanisms, which means that they are more vulnerable to small subtle perturbations that are distributed globally. Transferability is an important aspect of adversarial attacks, where perturbations designed for one model effectively impact another model, even if they differ architecturally. Research indicates that both Vision Transformers and CNNs exhibit this vulnerability, pointing out the need for robust defense strategies. For Example, Liao et al. elaborated on how adversarial examples can be transferred across different model architectures [2], while Xu et al. showed the effectiveness of various defense methods against these attacks, indicating that adversarial training can fortify models, including both CNNs and transformer-based architectures [3].
@@ 2.1 Adversarial Attacks and Watermarking in Deep Learning
Adversarial attacks pose significant threats to the integrity of deep learning models, particularly in sensitive applications such as medical imaging. The strategy of embedding watermarks has emerged as a method for securing digital assets against such attacks. Watermarking in deep learning used to authenticate the source and ownership of images, ensuring that any unauthorized alterations can be detected [4, 5]. In medical contexts, where data integrity is a must, these watermarking techniques not only protect the authenticity of the images but also deal with any potential security breaches [6]. The need for robust watermarking systems grows in importance as adversarial attacks become more sophisticated, underscoring their essential role in security strategies within the healthcare domain [7].
@@ 2.2 Traditional CNN-based Watermarking
Traditional watermarking techniques often leverage Convolutional Neural Networks (CNNs) to embed invisible marks into digital media. Approaches such as invisible embedding and frequency domain methods have been widely studied, where information is embedded in transformed image representations [8, 9]. However, these methods frequently encounter limitations regarding robustness and detection, as they may be susceptible to various forms of image manipulations, such as geometric transformations and noise, which impair watermark visibility and durability [10]. Despite addressing these issues to some extent, the inherent fragility of CNN-based schemes highlights the ongoing need for improvements in robustness [11].
@@ 2.3 Transformer-based Adversarial Watermarking
Recent advancements have begun to explore transformer architectures for watermarking within medical image processing. Research indicates that transformers provide enhanced feature extraction capabilities compared to traditional CNNs, facilitating the handling of diverse image characteristics encountered in medical datasets [12]. This shift not only leads to potentially better watermarking techniques but also offers a comparative advantage in versatility and robustness against various attack vectors [13]. Current studies suggest that the integration of transformer models can significantly enhance watermark reliability, although further empirical investigations are required to fully understand their efficacy in real-world applications [14, 15].
@@ 2.4 Transferability of Adversarial Watermarks
Transferability of adversarial watermarks between different models and datasets has emerged as a critical research area, particularly in the context of deep learning. Existing literature underscores the complexities involved in ensuring that watermarks remain intact across various architectures and data types [16]. However, there is a notable gap in studies specifically focusing on transformer watermark transferability in medical imaging domains, leaving a crucial area unexplored [17]. This lack of targeted research limits the understanding of how adversarial watermarks can function effectively in diverse model environments, particularly within healthcare applications where data migration and interoperability are prevalent challenges [18].
@@ 2.5 Medical Imaging Specific Context
In the medical imaging field, watermarking techniques are particularly sensitive to issues of security, ethical implications, and compliance with data protection regulations. Research has illustrated that certain watermarking methodologies designed for medical modalities (e.g., MRI, CT) face unique challenges, necessitating a tailored approach to ensure both invisibility and robustness [19, 20]. The integration of advanced watermarking solutions, such as using patient-specific data as a watermark, has been suggested as a means to bolster both security and ethical standards in healthcare [21]. Nevertheless, the application of these methods demands thorough validation to align with stringent medical standards concerning data integrity and patient privacy [17].
Despite advancements in adversarial watermarking, significant gaps persist in the literature. There is a pronounced deficiency in extensive studies addressing watermark transferability specifically within transformer models, limiting insights into their utility across varied healthcare applications. Moreover, insufficient robustness testing against advanced defensive techniques raises concerns about the practical viability of current methods in real-world scenarios [20]. Finally, limited practical experimentation using transformer architectures within prevalent medical imaging datasets underscores the need for focused research to bridge these gaps and enhance the security landscape in medical imaging systems.
@@ 4.1 Dataset and Preprocessing
We have used 5 classes of skin diseases, namely Atopic dermatitis, Eczema, Herpes, Nevus, and Melanoma. The dataset is used in a study conducted by Sadik et al. [22] where authors proposed an an automated system for skin disease recognition by implementing CNN architectures.
@@ 4.2 Results and Discussions
In this section, we present and analyze the empirical findings of our adversarial robustness experiments conducted on a skin disease image classification dataset. We evaluate three state-of-the-art deep learning architectures: Vision Transformer (ViT), ResNet-50, and VGG16, under both clean and adversarial conditions. Further, we analyze the impact of adversarial training on enhancing model robustness.
All three models were initially trained on clean images for comparing baseline classification accuracy in figure 5. The ViT and ResNet-50 models demonstrated strong performance, achieving accuracies of 94.4% and 94.2%, respectively, while VGG16 achieved 84.0%. These results confirm that transformer-based and deep CNN architectures are capable of effectively learning discriminative features for skin disease classification.
@@
To assess robustness, adversarial examples were crafted using the Projected Gradient Descent (PGD) method and this result has shown in figure 6. Under this attack, the classification accuracy of ViT dropped precipitously to 27.6%, while ResNet-50 and VGG16 degraded to 70.0% and 54.4%, respectively. This highlights a key vulnerability: ViT is significantly more sensitive to adversarial perturbations, likely due to its reliance on global self-attention mechanisms, which are easily disrupted by subtle pixel-level changes. In contrast, the localized feature extraction of CNNs offers more inherent resilience.
@@
The results confirm that adversarial training is highly effective, especially for ViT, which not only recovers robustness but also generalizes well across previously misclassified perturbations. This finding supports the notion that transformer models, while initially fragile, can become highly robust when trained with appropriate adversarial objectives.
In the context of medical imaging, where model predictions can directly impact diagnostic outcomes, adversarial robustness is paramount. The severe accuracy degradation observed in ViT prior to adversarial training underlines a critical risk in deploying unprotected transformer models in clinical settings. However, the successful recovery through adversarial training demonstrates that robust ViT models are viable and effective for real-world healthcare applications, provided they are sufficiently hardened.
@@ 5 Conclusions
In this paper, we have investigated the vulnerabilities of Vision Transformers (ViTs) to adversarial attacks for skin disease image classification task. The findings suggest that ViTs performed well on clean images, achieving maximum accuracy levels. It is also observed that the ViT model is vulnerable to adversarial perturbations as compared to traditional CNNs like ResNet-50 and VGG16. Proposed adversarial training showed success to significantly improve the robustness of ViTs against attacks from adversaries. However, adversarial training significantly improved ViT’s performance, recovering its adversarial accuracy from 27.6% to 90.0%, thereby demonstrating the model’s capacity to learn robust representations when properly regularized. Future work should focus on developing dynamic adversarial training strategies, exploring a broader range of preprocessing defenses, and evaluating the performance of ViTs on larger, more diverse datasets to ensure generalizability and resilience in real-world scenarios.

### v2

@@ 1 Introduction
Transfer based architectures, originally designed for natural language processing, have recently shown its promises in the field of image processing. Vision Transformers (ViTs), with their reliance on self-attention mechanisms, have demonstrated superior performance in tasks such as image recognition, segmentation, and multimodal learning [mao2022towards]. As the medical domain is known for the scarcity of data, transfer-based architectures will also play an important role in this field in terms of segmentation, anomaly detection, and eventually image recognition, i.e., diagnosis. Unlike traditional Convolutional Neural Networks (CNNs) that process localized spatial features, ViTs operate on global relationships across image patches, allowing them to capture long-range dependencies effectively. This enables them to outperform CNNs in specific scenarios, including subtle pattern identification in complex medical images. However, this global attention mechanism also exposes them to unique vulnerabilities, particularly in adversarial settings.
@@
ViT take images, in this case - medical images, and divide them up into patches of a fixed size, and then maps these into embeddings linearly. This patch-based representation lacks the sort of hierarchical feature extraction that even the simplest CNNs perform, where lower layers can recognize simple shapes and higher-order ones recognize much more complex features. In contrast, the attention mechanism in ViT considers all patches simultaneously for the computation of their relationships. Therefore, scaling perturbations that change the interrelationships of patches can be used to cause misrepresentation of attention maps with respect to the original structure of the image. ViTs also operate on random global attention mechanisms, which means that they are more vulnerable to small subtle perturbations that are distributed globally. Transferability is an important aspect of adversarial attacks, where perturbations designed for one model effectively impact another model, even if they differ architecturally. Research indicates that both Vision Transformers and CNNs exhibit this vulnerability, pointing out the need for robust defense strategies. For Example, Liao et al. elaborated on how adversarial examples can be transferred across different model architectures [liao2018defense], while Xu et al. showed the effectiveness of various defense methods against these attacks, indicating that adversarial training can fortify models, including both CNNs and transformer-based architectures [xu2021towards].
@@ 2.1 Adversarial Attacks and Watermarking in Deep Learning
Adversarial attacks pose significant threats to the integrity of deep learning models, particularly in sensitive applications such as medical imaging. The strategy of embedding watermarks has emerged as a method for securing digital assets against such attacks. Watermarking in deep learning used to authenticate the source and ownership of images, ensuring that any unauthorized alterations can be detected [Kaczyński, Mareen]. In medical contexts, where data integrity is a must, these watermarking techniques not only protect the authenticity of the images but also deal with any potential security breaches [LiDekai]. The need for robust watermarking systems grows in importance as adversarial attacks become more sophisticated, underscoring their essential role in security strategies within the healthcare domain [Clements].
@@ 2.2 Traditional CNN-based Watermarking
Traditional watermarking techniques often leverage Convolutional Neural Networks (CNNs) to embed invisible marks into digital media. Approaches such as invisible embedding and frequency domain methods have been widely studied, where information is embedded in transformed image representations [WuShiqiang, Dong]. However, these methods frequently encounter limitations regarding robustness and detection, as they may be susceptible to various forms of image manipulations, such as geometric transformations and noise, which impair watermark visibility and durability [Jabra]. Despite addressing these issues to some extent, the inherent fragility of CNN-based schemes highlights the ongoing need for improvements in robustness [ZhangWenxing].
@@ 2.3 Transformer-based Adversarial Watermarking
Recent advancements have begun to explore transformer architectures for watermarking within medical image processing. Research indicates that transformers provide enhanced feature extraction capabilities compared to traditional CNNs, facilitating the handling of diverse image characteristics encountered in medical datasets [LiuJing]. This shift not only leads to potentially better watermarking techniques but also offers a comparative advantage in versatility and robustness against various attack vectors [SinghHimanshu]. Current studies suggest that the integration of transformer models can significantly enhance watermark reliability, although further empirical investigations are required to fully understand their efficacy in real-world applications [Mahapatra, PuBangzheng].
@@ 2.4 Transferability of Adversarial Watermarks
Transferability of adversarial watermarks between different models and datasets has emerged as a critical research area, particularly in the context of deep learning. Existing literature underscores the complexities involved in ensuring that watermarks remain intact across various architectures and data types [Zhong]. However, there is a notable gap in studies specifically focusing on transformer watermark transferability in medical imaging domains, leaving a crucial area unexplored [LiuXinwei]. This lack of targeted research limits the understanding of how adversarial watermarks can function effectively in diverse model environments, particularly within healthcare applications where data migration and interoperability are prevalent challenges [review].
@@ 2.5 Medical Imaging Specific Context
In the medical imaging field, watermarking techniques are particularly sensitive to issues of security, ethical implications, and compliance with data protection regulations. Research has illustrated that certain watermarking methodologies designed for medical modalities (e.g., MRI, CT) face unique challenges, necessitating a tailored approach to ensure both invisibility and robustness [Nie, SinghKamred]. The integration of advanced watermarking solutions, such as using patient-specific data as a watermark, has been suggested as a means to bolster both security and ethical standards in healthcare [Zhang]. Nevertheless, the application of these methods demands thorough validation to align with stringent medical standards concerning data integrity and patient privacy [LiuXinwei].
Despite advancements in adversarial watermarking, significant gaps persist in the literature. There is a noticeable lack of extensive studies addressing watermark transferability specifically with transformer models. That limits their insights into their utility across varied healthcare applications. Moreover, insufficient robustness testing against advanced defensive techniques raises concerns about the practical viability of current methods in real-world scenarios [SinghKamred]. Finally, limited practical experimentation using transformer architectures within prevalent medical imaging datasets underscores the need for focused research to fill these gaps and improve security in medical imaging systems.
@@ 4.1 Dataset and Preprocessing
We have used 5 classes of skin diseases, namely Atopic dermatitis, Eczema, Herpes, Nevus, and Melanoma. The dataset is used in a study conducted by Sadik et al. [sadik2023depth] where authors proposed an an automated system for skin disease recognition by implementing CNN architectures.
@@ 4.2 Results and Discussions
This section presents the empirical findings of our adversarial robustness experiments, which were conducted on a skin disease image classification dataset. The effectiveness of three deep learning methods is justified with both clean and adversarial conditions. Also, the impact of adversarial training on improving the models’ robustness is analyzed.
At first, three models were trained on clean images for comparing baseline classification accuracy in figure 5. In this experiment, ViT and ResNet-50 delivered accuracies of 94.4% and 94.2%, respectively, while VGG16 achieved 84.0%. These results ensure the effectiveness of transformer-based and deep CNN architectures in learning discriminative features for skin disease classification.
@@
To verify the robustness of models, adversarial examples were crafted using the Projected Gradient Descent (PGD) method, and the results are shown in figure 6. During the attack, ViT performed very poorly with an accuracy of 27.6%, while ResNet-50 and VGG16 degraded to 70.0% and 54.4%, respectively. Therefore, it can be said that ViT is much more vulnerable to adversarial attacks due to the use of global self-attention that makes it sensitive to small pixel changes. But ResNet-50 and VGG16 are more resistant to these attacks because they focus on local features, making them more robust.
@@
The results confirm that adversarial training is highly effective for ViT. This method not only recovers robustness but also generalizes well across previously misclassified perturbations. This finding supports the notion that initially fragile ransformer models, can become highly robust when trained with appropriate adversarial objectives.
In medical imaging, where a model’s predictions can affect diagnoses, it’s very important for models to be robust against attacks. The sharp drop in ViT’s accuracy before adversarial training shows a major risk in using unprotected transformer models in healthcare. However, after adversarial training, ViT was able to recover and perform well, proving that these models can be effective and safe for real-world medical use if properly protected.
@@ 5 Conclusions
This paper presents the weaknesses of Vision Transformers (ViTs) to adversarial attacks in the context of skin disease image classification. According to this study, ViT can perform insanely well on clean images and achieved the most accurate results. In the case of adversarial perturbations, ViT performs poorly and provides less accuracy than traditional CNNs like ResNet-50 and VGG16. The proposed adversarial training method delivers exceptionally good performance using ViT against attacks from adversaries. After using adversarial training, ViT’s accuracy on adversarial images improved a lot from 27.6% to 90.0%. This demonstrates the model’s ability to learn robust representations with a properly regularized model.
Future research will concentrate on developing dynamic adversarial training strategies along with exploring different types of data preprocessing methods. Additionally, the effectiveness of ViT will be evaluated on larger and more diverse datasets to ensure generalizability and resilience in real-world scenarios.